Joe is 1.6m tall. His shadow is 2m long when he stands 3m from the base of a floodlight.
Show that the height of the flood light is equal to 4 m

Answer by mananth(16946)

Height of light pole and Joe are parallel
..
distance from foot of light = 3 feet
shadow = 2 feet
total = 5 feet
..
Joe is 1.6 feet.
Let height of light pole be x feet
by congruent triangle theorem
x/5 = 1.6 /2
x= 5*1.6/2
x= 4 the height of light